Given current and past global car annual production, estimate the quantity CO2 produced from operating these vehicles.
State all assumptions.

sorry I dont have that data ..the question is asked as it is I posted.
Answer #1

It is seen that, a typical passenger vehicle emits 4.6 metric tons Co2 in a year.

In past i.e in 2018, global car( typical passenger car) annual production is around 70 million.

In 2019, this is increased by one percent, i.e annual car ( typical passenger car) production is around 77 million.

In 2018, the quantity Co2 produced anually = 4.6* 70*10^6 = 322*10^6 metric ton in a year.

In 2019, the quantity of Co2 produce anually = 4.6* 77*10^6 = 354.2 *10^6 metric ton in a year.